OpenStudy (anonymous):

if sin(θ) = − 7/10 and sec(θ) >0, find the exact values of each of the trig functions how would you work this out? could someone please explain? cos(θ) = tan(θ) = cot(θ) = sec(θ) = csc(θ) =

OpenStudy (anonymous):

Solve for \(\cos(\theta)\) \[ \cos ^2(\theta )=1-\sin^2(\theta )=1-\left(-\frac{7}{10}\right)^2=\frac{51}{100}\\ \cos(\theta)=\pm \frac{\sqrt{51}}{10}\\ \Large \cos(\theta)= \frac{\sqrt{51}}{10}\\Since\\ \sec(\theta)>0 \]

OpenStudy (anonymous):

\[ \tan (\theta )=\frac{\sin (\theta )}{\cos (\theta )}=-\frac{7}{\frac{10 \sqrt{51}}{10}}=-\frac{7}{\sqrt{51}}\\ \cot(\theta)=-\frac{\sqrt{51}}7 \]

OpenStudy (anonymous):

\[ \Large \sec(\theta)= \frac{10}{\sqrt{51}}\\Since\\ \csc(\theta)=-\frac{10}7 \]

OpenStudy (anonymous):

thankyou so much for helping me again!